OpenStudy (anonymous):

what is my next step for sin 5 theta=1/2

OpenStudy (anonymous):

If you're trying to solve for θ, start by taking the inverse sin of both sides.

OpenStudy (dayakar):

\[\sin5\theta = \sin30\]

OpenStudy (dayakar):

\[since \sin30 = 1/2\]

OpenStudy (dayakar):

\[5\theta=30\]

OpenStudy (dayakar):

\[\theta=30/5\]

OpenStudy (dayakar):

\[\theta=6\]